OpenStudy (anonymous):

I don't get this. Help please! 1. -(6)^-1 A. 6 B. 1/6 C. -1/6 D. -6

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[\huge b^{-n}=\frac{1}{b^n}\] is what you need to kn ow

OpenStudy (misty1212):

that makes \[\huge -(6)^{-1}=-\frac{1}{6}\]

OpenStudy (misty1212):

in any case you can never go wrong with C, it is always C

OpenStudy (anonymous):

Ohh okay

OpenStudy (anonymous):

Thank you!

OpenStudy (misty1212):

yw this is what you need to "get" for these questions \[\huge b^{-n}=\frac{1}{b^n}\] don't never forget

OpenStudy (anonymous):

Okay. Thanks a lot :)

OpenStudy (misty1212):

yw